# YITP-26-12, RESCEU-4/26, IPMU26-0003

### v1

@@ Abstract
This letter proposes a new avenue for understanding the cosmological singularity. The standard cosmological model contains a generic initial singularity usually referred to as the big bang. Herein, we present a novel idea to extend the description of our Universe beyond this limit. The proposal relies on rewriting physics in a purely Riemannian, i.e. locally Euclidean, 4-dimensional space and the emergence of Lorentzian patches owing to the interaction of all matter fields to a clock field that is responsible for a signature change. If our universe is contained within one of these patches, the initial singularity is replaced by a smooth boundary on which the signature of the physical metric flips. In this letter, we first define the model and draw the necessary conditions on its arbitrary functions for solutions to exist. Next, we prove the existence of solutions that lead to an emergent universe with a primordial (almost) de Sitter phase. To finish, we discuss the consequences of this construction for the universe on scales much larger than our observable universe: a large “Euclidean sea” in which Lorentzian islands locally emerge and host an expanding universe potentially similar to ours. While speculative, this scenario has specific features that can be tested, and the present letter sets the basis for further phenomenological investigations.
@@ I Introduction
The existence of the primordial singularity of the standard cosmological model has long been identified as a signal of the domain of validity of general relativity to describe gravitation, starting perhaps with Lemaître’s early insight Lemaitre:1931zzb . Many classical solutions to avoid the singularity within general relativity have been proposed, such as the Eddington-Lemaître model Eddington:1930zz ; Lemaitre:1931zza , bouncing solutions that both require a non-vanishing spatial curvature, i.e. an effective fluid with equation of state $P=-\rho/3$ with positive or negative energy density according to the sign of the spatial curvature, oscillating universes Tolman:1931fei ; Lemaitre:1933gd ; Ellis:2002we or even a continuous creation of matter Hoyle:1948zz . Indeed those solutions are either not compatible with the actual measurements of the cosmological parameters or do not avoid the singularity.
@@
The occurrence of singularities in general relativity has been related to energy conditions Hawking:1966vg ; Hawking:1966sx ; Hawking:1973uf and those early classical solutions have all seen counterparts in the developments of quantum gravity phenomenology, including bouncing universes thanks to duality Veneziano:1991ek ; Brandenberger:1988aj , the dynamics of the dilaton Gasperini:1992em , branes Steinhardt:2002ih ; Steinhardt:2001st in string theory, or the the quantum pressure Rovelli:1997yv ; Bojowald:2005epg that appears in loop quantum gravity. On the side of quantum cosmology, many solutions, invoking the non-boundary proposal Hartle:1983ai , tunneling geometries Vilenkin:1982de , or phase transitions in spin networks models Mielczarek:2012pf inspired by signature change in meta-materials Smolyaninov:2010tpq , have been considered to describe the emergence of our universe.
@@
It is worth noting that one can find solutions of the Einstein equations for a metric with Lorentzian as well as Euclidean signatures. It was actually conjectured by Sakharov Sakharov:1984csx , inspired by Ref. Vilenkin:1982de , that “there may exist states of the physical continuum which include regions with different signatures of the metric and that the observed Universe and an infinite number of other universes arose as a result of quantum processes with a change in the signature of the metric”. This led several authors Ellis:1991sp ; Ellis:1991st ; Hayward:1992zp ; Dray:1996cw ; Dray:1996dc to study the junction conditions between such spacetime patches with different signatures at the classical level, which require these regions to be glued smoothly on their boundaries with an either smooth Ellis:1991sp or discontinuous Ellis:1991st ; Hayward:1992zp sign shift. However, these studies assume a profile of sign changing $g_{00}$ and do not construct a full model of time emergence based on a consistent field theory as proposed in Refs. Mukohyama:2013ew ; Kehayias:2014uta —see also earlier work in the context of Nordström gravity Girelli:2008qp .
@@
Hence, following our previous investigations Mukohyama:2013ew ; Kehayias:2014uta ; Mukoyama:2013gqu ; Mukohyama:2013gra ; Uzan-CUP ; Feng:2023klt ; Feng:2025xsi on a field theory based on a purely Euclidean (hereafter, by “Euclidean” we mean “Riemannian”, i.e. “locally Euclidean”) space, we propose in this letter a new road to understanding the big bang singularity. This theoretical approach relies on the hypothesis that (i) at the microscopic level the metric is Riemannian, i.e., that all matter fields are defined in a 4-dimensional Euclidean space with metric of signature $(+,+,+,+)$,
@@
$g_{\mu\nu}=g^{\rm E}_{\mu\nu}-\frac{\partial_{\mu}\phi\partial_{\nu}\phi}{M^{4}}\,$ (2)
@@
where $M$ is a mass scale. Refs. Mukohyama:2013ew ; Kehayias:2014uta ; Mukoyama:2013gqu ; Mukohyama:2013gra ; Uzan-CUP ; Feng:2023klt ; Feng:2025xsi demonstrated that if the gradient of the clock field develops a constant non-vanishing vacuum expectation value on a patch ${\cal M}_{0}$, then within this patch the effective metric can exhibit a (,+,+,+)(-,+,+,+) signature. Furthermore, it was shown that the whole action of the standard model of particle physics in the patch ${\cal M}_{0}$ can be obtained while gravity is described by a more general covariant Galileon theory. While in the standard approach of any field theory, the Lorentzian metric structure allows one to implement the relativistic notion of causality and to define a notion of time dimension, it is now seen as an emergent property. Hence, in this description, there is no dynamics nor signature flip across some hypersurface; instead, all the fields develop a Lorentzian dynamics in some patches of space because they minimally couple to an effective metric that depends on the local value of the clock field.
@@
This insight offers a new view on the notion of time as an emergent property. From a phenomenological point of view, this shall affect the construction of our cosmological model since the whole dynamical observable universe with a well-defined time arrow can exist only within the patch ${\cal M}_{0}$. On the boundaries $\Sigma_{0}$ of ${\cal M}_{0}$, “time freezes” since the space becomes Euclidean but without necessary passing through a singularity. Hence, one is led to the idea that if our universe is contained in such a patch ${\cal M}_{0}$ then the cosmological singularity of the big bang is replaced by the boundary $\Sigma_{0}$ of this patch.
@@
To investigate this idea further and make it more concrete, we shall construct a 4-dimensional Euclidean space with planar symmetry such that both $g^{\rm E}_{\mu\nu}$ and $\phi$ depend only on one coordinate, $z$, so that the $\{z={\rm const.}\}$ hypersurfaces are homogeneous and isotropic, ensuring that the emergent universe [^1] To avoid ambiguities, we shall use the term universe for the Lorentzian regions for which we can use the words “time” and “space” in a well-defined way. satisfies the Copernican principle. Indeed, if a shift of signature (2) occurs within ${\cal M}_{0}$ the $z$ direction will play the role of the arrow of time since $\{z={\rm const.}\}$ will be $\{\phi={\rm const.}\}$ and thus $\{t={\rm const.}\}$. We further assume that the effective metric has the signature (,+,+,+)(-,+,+,+) in the asymptotic regions. The whole global space structure is as described in Fig. 1, that is, two “mirror” universes ${\cal M}_{0+}$ and ${\cal M}_{0-}$ with signature (,+,+,+)(-,+,+,+) match a central Euclidean region with no time. Each of the two Lorentzian zones can be thought as a universe, with time flowing in opposite directions. What the observers in each emergent universe will identify as the big bang is simply the boundary $\Sigma_{0}$ on which the signature of the metric flips.
@@
Figure 1: A big bang that never happened. Our universe can be thought as an emergent Lorentzian patch within a 4-dimensional purely Riemannian space. The “big bang” can be understood as the hypersurface $\Sigma_{0}$ on which the signature of the metric flips. In the simplest version, the geometry and field configurations are assumed to (i) depend only on one space coordinate, $z$, so that the emergent spacetimes ${\cal M}_{0\pm}$ satisfy the Copernican principle and (ii) exhibit the (,+,+,+)(-,+,+,+) signature for the effective metric in the two asymptotic regions.
@@ II.1 Definition of the theory
We decompose the Euclidean metric $g^{\rm E}_{\mu\nu}$ in terms of the lapse $N_{E}$ and the metric $\gamma_{ij}$ on the 3-dimensional hypersurface normal to a coordinate $z$ as
@@
Indeed, $h=\ln a$ so that $\dot{h}={\dot{a}}/{a}$ with a dot denoting a derivative with respect to $z$, and for latter convenience we define
@@
so that the configuration up to a constant shift of $\phi$ is characterized by the 3 functions $N_{\rm E}$, $h$ and $\psi$ of the variable $z$. $N_{\rm E}$ can always be set to unity thanks to a redefinition of $z$ so that we only have effectively 2 functions to determine the model.
Following Ref. Mukohyama:2013ew – see its Eq. (3.6) – we consider the action $S=\int{\cal L}\mathrm{d}^{4}x$ with Lagrangian
@@
with $\nabla^{2}_{\rm E}$ the covariant derivative associated to $g^{\mu\nu}_{\rm E}$ and $\nabla^{2}_{\rm E}\phi\equiv g^{\mu\nu}_{\rm E}\nabla_{\mu}\nabla_{\nu}\phi=(\dot{\psi}+3\dot{h}\psi)/N_{\rm E}$. The action involves only the two free functions, $\cal K$} and $G_{4}$, that depend only on $X_{\rm E}$ defined as
@@
Indeed only 2 of the 3 equations (9-11) are independent thanks to the Bianchi identity. Hence, for the sake of simplicity, we only use the first two (9-10) in our analysis. Defining
@@ II.2 Necessary condition for the existence of solutions
Suppose that $\psi(z=z_{0})=0$ and that $j(z=z_{0})$ is finite and non-zero. Then $J_{0}=0$ and thus $j(z)\psi(z)=0$ for all $z$. Also, by continuity $j(z)$ is finite and non-zero in an open neighborhood of $z=z_{0}$. These two facts imply that $\psi(z)=0$ in the open neighborhood, meaning in particular that $\psi$ and its all-order derivatives vanish at $z=z_{0}$. Since $\phi(z)$ satisfies a second-order ordinary differential equation and $\psi(z)$ is its first derivative, it is concluded that $\psi(z)=0$ for all $z$. Therefore, $\psi$ cannot cross zero unless $j$ also vanishes there.
To evade this no-go result, we can for example consider clock-field configurations with $Z_{2}$ antisymmetry”, i.e.,
@@
with $\phi_{0}$ an arbitrary constant, and metric configurations with $Z_{2}$ symmetry”,
@@
In such a case, the existence of a non-trivial solution with $j\to 0$ ($z\to\pm\infty$) is not forbidden by the no-go and could be found by tuning $\psi(z=0)$. If $H_{\rm E}(z)>0$ for $z>0$ then this is basically what we could call a Euclidean bouncing universe. The shift charge density $j\psi$ will automatically approach zero as $z\to\pm\infty$ if $\psi(z=0)$ is in the basin of the attractor.
@@
The previous argument teaches us that one has to set boundary conditions such that $\psi$ tends toward either zero or a constant that makes $j$ vanish as $z\to\pm\infty$, and that they never cross $\psi=0$. In order to have two external zones with Lorentz signature and a de Sitter geometry at $z=\pm\infty$, one needs to ensure that Eq. (20) holds and that $H_{\rm E}$ approaches a positive constant as $z\to\pm\infty$ .
@@ II.3 Cosmological singularity
It follows that while the scale factor remains on $\Sigma_{0}:\{\psi^{2}=M^{4}\}$, $H$ will be singular on $\Sigma_{0}$ if $H_{\rm E}$ remains non-vanishing.
The singular behavior in $H$ should not be surprising, as the effective metric becomes degenerate on the hypersurface $\Sigma_{0}$. It has been pointed out in the literature that while a degenerate metric may correspond to coordinate singularities, a degenerate metric on a surface of codimension one may signal the presence of a distributional or thin-shell Feng:2023vqp singularity (cf. also Refs. Ellis:1991sp ; Ellis:1991st ; Hayward:1992zp ; Dray:1996cw ; Dray:1996dc ), and other types of curvature singularities (as in the example presented in Ref. Feng:2023klt ). In our scenario the hypersurface on which the effective metric becomes degenerate is completely regular from the viewpoint of the Euclidean theory and corresponds to the cosmological singularity from the viewpoint of the emergent spacetime.
@@ II.4 General behavior
Up to sign changes, $H_{\rm E}$ and $a$ are thus expressed algebraically as functions of $X_{\rm E}$. One can see that these equations implicitly specify the dependence of the system on $z$ through the relation $(\ln a^{3})^{.}=3H_{\rm E}$. One can obtain a differential equation for $\psi$ by differentiating $a^{3}(X_{\rm E})$ with respect to $\psi$, employing the chain rule $\dot{F}=\dot{X}_{\rm E}F^{\prime}=2\psi\dot{\psi}F^{\prime}$ to obtain:
@@
The expression for the derivative $\dot{H}_{\rm E}/N_{\rm E}$ is obtained from the chain rule and the expression ${\dot{\psi}}/{N_{\rm E}}$. Observe that the r.h.s. of each expression in Eq. (27) may be expressed in terms of $\psi=\sqrt{X_{\rm E}}$ using the expressions in Eq. (26). This allows us to extract the general behavior of the system from the functional forms of $\mathcal{K}(X_{\rm E})$ and $\mathcal{G}(X_{\rm E})$ without solving the differential equations. Of course, if the r.h.s. of Eq. (26) is expressed explicitly in terms of $\psi$, one may also read it as a first-order differential equation for $\psi(z)$; given the solution $\psi(z)$, the algebraic expressions (26) supply the $z$ dependence (up to sign changes) for $H_{\rm E}$ and $a$.

### v2

@@ Abstract
This paper proposes a new avenue for understanding the cosmological singularity. The standard cosmological model contains a generic initial singularity usually referred to as the big bang. Herein, we present a novel idea to extend the description of our Universe beyond this limit. The proposal relies on rewriting physics in a purely Riemannian, i.e. locally Euclidean, four-dimensional space and the emergence of Lorentzian patches owing to the interaction of all matter fields to a clock field that is responsible for a signature change. If our Universe is contained within one of these patches, the initial singularity is replaced by a smooth boundary on which the signature of the physical metric flips. In this paper, we first define the model and draw the necessary conditions on its arbitrary functions for solutions to exist. Next, we prove the existence of solutions that lead to an emergent universe with a primordial (almost) de Sitter phase. To finish, we discuss the consequences of this construction for the universe on scales much larger than our observable Universe: a large “Euclidean sea” in which Lorentzian islands locally emerge and host an expanding universe potentially similar to ours. While speculative, this scenario has specific features that can be tested, and the present paper sets the basis for further phenomenological investigations.
@@ I Introduction
The existence of the primordial singularity of the standard cosmological model has long been identified as a signal of the domain of validity of general relativity to describe gravitation, starting perhaps with Lemaître’s early insight Lemaitre:1931zzb . Many classical solutions to avoid the singularity within general relativity have been proposed, such as the Eddington-Lemaître model Eddington:1930zz ; Lemaitre:1931zza , bouncing solutions that both require a nonvanishing spatial curvature, i.e. an effective fluid with equation of state $P=-\rho/3$ with positive or negative energy density according to the sign of the spatial curvature, oscillating universes Tolman:1931fei ; Lemaitre:1933gd ; Ellis:2002we or even a continuous creation of matter Hoyle:1948zz . Indeed those solutions either are not compatible with the actual measurements of the cosmological parameters or do not avoid the singularity.
@@
The occurrence of singularities in general relativity has been related to energy conditions Hawking:1966vg ; Hawking:1966sx ; Hawking:1973uf and those early classical solutions have all seen counterparts in the developments of quantum gravity phenomenology, including bouncing universes thanks to duality Veneziano:1991ek ; Brandenberger:1988aj , the dynamics of the dilaton Gasperini:1992em , branes Steinhardt:2002ih ; Steinhardt:2001st in string theory, or the the quantum pressure Rovelli:1997yv ; Bojowald:2005epg that appears in loop quantum gravity. On the side of quantum cosmology, many solutions, invoking the no-boundary proposal Hartle:1983ai , tunneling geometries Vilenkin:1982de , or phase transitions in spin network models Mielczarek:2012pf inspired by signature change in metamaterials Smolyaninov:2010tpq , have been considered to describe the emergence of our Universe.
@@
It is worth noting that one can find solutions of the Einstein equations for a metric with Lorentzian as well as Euclidean signatures. It was actually conjectured by Sakharov Sakharov:1984csx , inspired by Ref. Vilenkin:1982de , that “there may exist states of the physical continuum which include regions with different signatures of the metric and that the observed Universe and an infinite number of other universes arose as a result of quantum processes with a change in the signature of the metric.” This led several authors Ellis:1991sp ; Ellis:1991st ; Hayward:1992zp ; Dray:1996cw ; Dray:1996dc to study the junction conditions between such spacetime patches with different signatures at the classical level, which require these regions to be glued smoothly on their boundaries with an either smooth Ellis:1991sp or discontinuous Ellis:1991st ; Hayward:1992zp sign shift. However, these studies assume a profile of sign changing $g_{00}$ and do not construct a full model of time emergence based on a consistent field theory as proposed in Refs. Mukohyama:2013ew ; Kehayias:2014uta —see also earlier work in the context of Nordström gravity Girelli:2008qp .
@@
Hence, following our previous investigations Mukohyama:2013ew ; Kehayias:2014uta ; Mukoyama:2013gqu ; Mukohyama:2013gra ; Uzan-CUP ; Feng:2023klt ; Feng:2025xsi on a field theory based on a purely Euclidean (hereafter, by “Euclidean” we mean “Riemannian,” i.e. “locally Euclidean”) space, we propose in this paper a new road to understanding the big bang singularity. This theoretical approach relies on the hypothesis that (i) at the microscopic level the metric is Riemannian, i.e., that all matter fields are defined in a four-dimensional Euclidean space with metric of signature $(+,+,+,+)$,
@@
$g_{\mu\nu}=g^{\rm E}_{\mu\nu}-\frac{\partial_{\mu}\phi\partial_{\nu}\phi}{M^{4}}\,$, (2)
@@
where $M$ is a mass scale. References Mukohyama:2013ew ; Kehayias:2014uta ; Mukoyama:2013gqu ; Mukohyama:2013gra ; Uzan-CUP ; Feng:2023klt ; Feng:2025xsi demonstrated that if the gradient of the clock field develops a constant nonvanishing vacuum expectation value on a patch ${\cal M}_{0}$, then within this patch the effective metric can exhibit a (,+,+,+)(-,+,+,+) signature. Furthermore, it was shown that the whole action of the standard model of particle physics in the patch ${\cal M}_{0}$ can be obtained while gravity is described by a more general covariant Galileon theory. While in the standard approach of any field theory, the Lorentzian metric structure allows one to implement the relativistic notion of causality and to define a notion of time dimension, it is now seen as an emergent property. Hence, in this description, there is no dynamics nor signature flip across some hypersurface; instead, all the fields develop a Lorentzian dynamics in some patches of space because they minimally couple to an effective metric that depends on the local value of the clock field.
@@
This insight offers a new view on the notion of time as an emergent property. From a phenomenological point of view, this shall affect the construction of our cosmological model since the whole dynamical observable Universe with a well-defined time arrow can exist only within the patch ${\cal M}_{0}$. On the boundaries $\Sigma_{0}$ of ${\cal M}_{0}$, “time freezes” since the space becomes Euclidean but without necessary passing through a singularity. Hence, one is led to the idea that if our Universe is contained in such a patch ${\cal M}_{0}$, then the cosmological singularity of the big bang is replaced by the boundary $\Sigma_{0}$ of this patch.
@@
To investigate this idea further and make it more concrete, we shall construct a four-dimensional Euclidean space with planar symmetry such that both $g^{\rm E}_{\mu\nu}$ and $\phi$ depend only on one coordinate, $z$, so that the $\{z={\rm const.}\}$ hypersurfaces are homogeneous and isotropic, ensuring that the emergent universe [^1] To avoid ambiguities, we shall use the term universe for the Lorentzian regions for which we can use the words “time” and “space” in a well-defined way. satisfies the Copernican principle. Indeed, if a shift of signature (2) occurs within ${\cal M}_{0}$ the $z$ direction will play the role of the arrow of time since $\{z={\rm const.}\}$ will be $\{\phi={\rm const.}\}$ and thus $\{t={\rm const.}\}$. We further assume that the effective metric has the signature (,+,+,+)(-,+,+,+) in the asymptotic regions. The whole global space structure is as described in Fig. 1; that is, two “mirror” universes ${\cal M}_{0+}$ and ${\cal M}_{0-}$ with signature (,+,+,+)(-,+,+,+) match a central Euclidean region with no time. Each of the two Lorentzian zones can be thought as a universe, with time flowing in opposite directions. What the observers in each emergent universe will identify as the big bang is simply the boundary $\Sigma_{0}$ on which the signature of the metric flips.
@@
Figure 1: A big bang that never happened. Our Universe can be thought as an emergent Lorentzian patch within a four-dimensional purely Riemannian space. The “big bang” can be understood as the hypersurface $\Sigma_{0}$ on which the signature of the metric flips. In the simplest version, the geometry and field configurations are assumed to (i) depend only on one space coordinate, $z$, so that the emergent spacetimes ${\cal M}_{0\pm}$ satisfy the Copernican principle and (ii) exhibit the (,+,+,+)(-,+,+,+) signature for the effective metric in the two asymptotic regions.
@@ II.1 Definition of the theory
We decompose the Euclidean metric $g^{\rm E}_{\mu\nu}$ in terms of the lapse $N_{E}$ and the metric $\gamma_{ij}$ on the three-dimensional hypersurface normal to a coordinate $z$ as
@@
Indeed, $h=\ln a$ so that $\dot{h}={\dot{a}}/{a}$ with a dot denoting a derivative with respect to $z$, and for later convenience we define
@@
so that the configuration up to a constant shift of $\phi$ is characterized by the three functions $N_{\rm E}$, $h$ and $\psi$ of the variable $z$. $N_{\rm E}$ can always be set to unity thanks to a redefinition of $z$ so that we only have effectively two functions to determine the model.
[…]
@@
with $\nabla^{\rm E}_{\mu}$ the covariant derivative associated to $g^{\mu\nu}_{\rm E}$, $\nabla^{2}_{\rm E}\phi\equiv g^{\mu\nu}_{\rm E}\nabla_{\mu}\nabla_{\nu}\phi=(\dot{\psi}+3\dot{h}\psi)/N_{\rm E}$, and $\left(\nabla_{\mu}\nabla_{\nu}\phi\right)^{2}\equiv g^{\mu\sigma}_{\rm E}g^{\nu\tau}_{\rm E}\nabla_{\mu}\nabla_{\nu}\phi\nabla_{\sigma}\nabla_{\tau}\phi$. The action involves only the two free functions, $\cal K$} and $G_{4}$, that depend only on $X_{\rm E}$ defined as
@@
Indeed only two of the three equations (9-11) are independent thanks to the Bianchi identity. Hence, for the sake of simplicity, we only use the first two (9 and 10) in our analysis. Defining
@@ II.2 Necessary condition for the existence of solutions
Suppose that $\psi(z=z_{0})=0$ and that $j(z=z_{0})$ is finite and nonzero. Then $J_{0}=0$ and thus $j(z)\psi(z)=0$ for all $z$. Also, by continuity $j(z)$ is finite and nonzero in an open neighborhood of $z=z_{0}$. These two facts imply that $\psi(z)=0$ in the open neighborhood, meaning in particular that $\psi$ and its all-order derivatives vanish at $z=z_{0}$. Since $\phi(z)$ satisfies a second-order ordinary differential equation and $\psi(z)$ is its first derivative, it is concluded that $\psi(z)=0$ for all $z$. Therefore, $\psi$ cannot cross zero unless $j$ also vanishes there.
To evade this no-go result, we can for example consider clock-field configurations with $Z_{2}$ antisymmetry,” i.e.,
@@
with $\phi_{0}$ an arbitrary constant, and metric configurations with $Z_{2}$ symmetry,”
@@
In such a case, the existence of a nontrivial solution with $j\to 0$ ($z\to\pm\infty$) is not forbidden by the no-go and could be found by tuning $\psi(z=0)$. If $H_{\rm E}(z)>0$ for $z>0$ then this is basically what we could call a Euclidean bouncing universe. The shift charge density $j\psi$ will automatically approach zero as $z\to\pm\infty$ if $\psi(z=0)$ is in the basin of the attractor.
@@
The previous argument teaches us that one has to set boundary conditions such that $\psi$ tends toward either zero or a constant that makes $j$ vanish as $z\to\pm\infty$ and that they never cross $\psi=0$. In order to have two external zones with Lorentz signature and a de Sitter geometry at $z=\pm\infty$, one needs to ensure that Eq. (20) holds and that $H_{\rm E}$ approaches a positive constant as $z\to\pm\infty$ .
@@ II.3 Cosmological singularity
It follows that while the scale factor remains on $\Sigma_{0}:\{\psi^{2}=M^{4}\}$, $H$ will be singular on $\Sigma_{0}$ if $H_{\rm E}$ remains nonvanishing.
The singular behavior in $H$ should not be surprising, as the effective metric becomes degenerate on the hypersurface $\Sigma_{0}$. It has been pointed out in the literature that while a degenerate metric may correspond to coordinate singularities, a degenerate metric on a surface of codimension one may signal the presence of a distributional or thin-shell Feng:2023vqp singularity (cf. also Refs. Ellis:1991sp ; Ellis:1991st ; Hayward:1992zp ; Dray:1996cw ; Dray:1996dc ) and other types of curvature singularities (as in the example presented in Ref. Feng:2023klt ). In our scenario the hypersurface on which the effective metric becomes degenerate is completely regular from the viewpoint of the Euclidean theory and corresponds to the cosmological singularity from the viewpoint of the emergent spacetime.
@@ II.4 General behavior
Up to sign changes, $H_{\rm E}$ and $a$ are thus expressed algebraically as functions of $X_{\rm E}$. One can see that these equations implicitly specify the dependence of the system on $z$ through the relation $(\ln a^{3})^{.}=3H_{\rm E}$. One can obtain a differential equation for $\psi$ by differentiating $a^{3}(X_{\rm E})$ with respect to $\psi$, employing the chain rule $\dot{F}=\dot{X}_{\rm E}F^{\prime}=2\psi\dot{\psi}F^{\prime}$ to obtain
@@
The expression for the derivative $\dot{H}_{\rm E}/N_{\rm E}$ is obtained from the chain rule and the expression ${\dot{\psi}}/{N_{\rm E}}$. Observe that the rhs of each expression in Eq. (27) may be expressed in terms of $\psi=\sqrt{X_{\rm E}}$ using the expressions in Eq. (26). This allows us to extract the general behavior of the system from the functional forms of $\mathcal{K}(X_{\rm E})$ and $\mathcal{G}(X_{\rm E})$ without solving the differential equations. Of course, if the rhs of Eq. (26) is expressed explicitly in terms of $\psi$, one may also read it as a first-order differential equation for $\psi(z)$; given the solution $\psi(z)$, the algebraic expressions (26) supply the $z$ dependence (up to sign changes) for $H_{\rm E}$ and $a$.